OpenStudy (ksaimouli):

a particle is moving with an initial velocity of +112m/s at t=0. The particle has no acceleration until t=3s, after which it has an acceleration of -4.om/s^2. what is the velocity at t=16.0s?

OpenStudy (anonymous):

Hint: use v=u+at

OpenStudy (anonymous):

u=112m/s ...... t = 16-3=13s and a=-4m/s^2

OpenStudy (shane_b):

@ksaimouli: This was answered in the Physics group too...please don't cross-post :)